OpenStudy (anonymous):

simplify √0.0016 0.04 4 0.004 0.4

OpenStudy (studygurl14):

What is 0.0016 in fraction form?

OpenStudy (anonymous):

16/10000 @Studygurl14

OpenStudy (studygurl14):

um, yes. But what's that simplified?

OpenStudy (anonymous):

i dont quite know :/

OpenStudy (radar):

Or you could express it as\[\sqrt{16}\sqrt{10^{-4}}\]

OpenStudy (anonymous):

hmm i'm really confused

OpenStudy (studygurl14):

@radar hmm...I didn't think of that

OpenStudy (studygurl14):

@radar 's method is better @bayan143

OpenStudy (radar):

either way will work. The method I am showing requires the student to know the rules for exponents. And scientific notation. I don't know where bayan143 is at in those subjects. How about it bayan143?

OpenStudy (radar):

Then do the StudyGurl14 method expressing .0016 as a fraction.

OpenStudy (anonymous):

ok :/

OpenStudy (radar):

.1 = 1/10 .01 = 1/100 .001=1/100 etc Now do that with .0016

OpenStudy (radar):

.16 = 16/100 .016 = 16/1000 what would .0016 look like as a fraction?

OpenStudy (radar):

When you get back, .0016 is the same as 16/10000 now you are going to solve for